UT 20161025 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt
Observer AO: DMiller
Telescope Operator: GBechetti

Plan for the night

Clouded out, hence until 10pm:

- Closed-dome AFC/PFC-tests

- Stability G040

Summary

- AFC works for LUCI2, but not for LUCI1

- no obvious issue with G040

Details

All times are UT
Files are luci1.20161025.NNNN.fits

00:45 Sunset, overcast, T=8deg, hum=60%

AFC/PFC/noFC-tests

01:31 N30_noFCtest_twin, set RA to 20h45min, LST is 20h30min

L1: #0139-0155

L2. #0025-0041

Stopped at 01:42, geno moves telescope to horizon due to heavy rain

02:06 Genos moves telescope back to zenith

02:12 N30_noFCtest_twin, set RA to 21h24min, LST is 21h08min

L1: #0156-0185

L2: #0043-0072

02:50 N30_PFCtest_twin, set RA to 22h04min, LST is 21h47min

L1: #0186-0215

L2: #0073-0102

02:38 N30_AFCtest_twin, set RA to 22h45min, LST is 22h25min

L1: #0216-0233

L2: #0103-0120

Did verification after alignment, needed to be redone in L2, both looked OK than

seems to work right away, correction <= 5steps n x,y on both sides

Luci1-AFC went bananas during the2nd image (again towards NNE as we say last night on sky)

Is there wrong scaling-factor for either (or both) comp.mirrors been applied?

LUCI2-AFC did its job

Darks:

L1: SRR, 120sec, #0234, LIR, 60sec, #0235

L2: SRR, 120sec, #0121, LIR, 60sec, #0122

Tests G040

N3.75, 0.25" slit, G040, Ar

Voltage set via Hiramo

123: testimage

124: 0V, line@1160
125: -0.25V, line@963
126: 0.25V, line@1356
127: 0V, line@1160
128: -0.25, line@962
129: 0.25, line@1356
130: -0.25, line@962
131: 0V, line @1160

move to G210, than back to G040

132: 0V, line@1172
133: -0.25, line@970

move to G200, than back to G040

134: 0.25V, line@1365

move to mirror, than back to G040

135: 0V, line@1166

move to g210, g200, mirror, g040

136: -0.25V, line@966
137: -2v, line@1491,
138. 2v, out of range no line
139: 0v, line @1170

move to g200, back to g040

139: -2v, line@1495

move to 210, back to g040

140: -0.25v, line@966
141: 0.25, line@1365
142: 0v, line@1165

seems all OK

end of tests

-- %USERSIG{JochenHeidt - 2016-10-25}%
Topic revision: r2 - 26 Oct 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback